OpenStudy (anonymous):

When solving the system of equations by graphing, what is the solution of 3x + 2y = 2 and 2x – y = 6? A. (–2, 2) B. (2, –2) C. (–2, –2) D. (2, 2)

OpenStudy (anonymous):

Do you have to use graphing?

OpenStudy (anonymous):

nooooo

OpenStudy (anonymous):

Which method do you want to use solving? Addition, Subtraction, or Substitution?

OpenStudy (anonymous):

addition

OpenStudy (anonymous):

This is the same problem ... lol

OpenStudy (anonymous):

oh so wa the answer

OpenStudy (anonymous):

Let me do it first.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You want to do addition right? You have 3x+2y=2 2x-y=6 You want to eliminate a variable by addition Multiply the second equation by 2 Now it becomes 4x-2y=12 3x+2y=2 4x-2y=12 Add the equations 7x=14 x=2 You want to find y now so plug 2 back into one of the equations. Let's use the first one 3(2) +2y=2 6+2y=2 2y=-4 y=-2 So your answer is (2,-2)

OpenStudy (anonymous):

hich of the following shows the solution for this system of linear equations? y = 2x – 1 2003-11-01-00-00_files/i0030000.jpg A. 2003-11-01-00-00_files/i0030001.jpg B. 2003-11-01-00-00_files/i0030002.jpg C. 2003-11-01-00-00_files/i0030003.jpg D. 2003-11-01-00-00_files/i0030004.jpg

OpenStudy (anonymous):

Please post this as a new question, rules are new question new post.

OpenStudy (anonymous):

ok